OpenStudy (anonymous):

The length of a rectangle is 2 cm less than 7 times the width . The perimeter is 60 . Find the width and length .

OpenStudy (jdoe0001):

7 times "something" => 7(something) 2 less than "something" => 2 - (something) 2 less than 7 times "something" => 2-7(something) perimeter is just the SUM OF THE SIDES a rectangle has 4 SIDES, 2 equal lengths, 2 equal width -------------------------------------------------------- if the "width" is say "w", the length will be 2-7w W = w L = 2-7w perimiter is W + W + L + L => 2W + 2W = > 2w + 2(2-7w) ------------------------------------------- we know the perimeter is 60, thus \(\bf perimeter = W + W + L + L => 2W + 2W = > \color{blue}{2w + 2(2-7w)}\\\qquad \\ 60 = 2w + \color{blue}{2(2-7w)}\) solve for "w"